OpenStudy (anonymous):

sin theta= -3/5 cos theta= 4/5

OpenStudy (tkhunny):

Have you considered drawing a right triangle? Of course, this is only guessing at what you might need, since you didn't actually post a question. You also showed none of your own work.

OpenStudy (anonymous):

i need to find the four remaining trigonometric functions

OpenStudy (tkhunny):

Well, then do it. tan(x) = sin(x)/cos(x) sec(x) = 1/cos(x) csc(x) = 1/sin(x) Only one more to go. Let's see it.